# Inverse k-visibility for RSSI-based Indoor Geometric Mapping

### v1

@@ Abstract
In recent years, the increased availability of WiFi in indoor environments has gained an interest in the robotics community to leverage WiFi signals for enhancing indoor SLAM (Simultaneous Localization and Mapping) systems. SLAM technology is widely used, especially for the navigation and control of autonomous robots. This paper discusses various works in developing WiFi-based localization and challenges in achieving high-accuracy geometric maps. This paper introduces the concept of inverse k-visibility developed from the k-visibility algorithm to identify the free space in an unknown environment for planning, navigation, and obstacle avoidance. Comprehensive experiments, including those utilizing single and multiple RSSI signals, were conducted in both simulated and real-world environments to demonstrate the robustness of the proposed algorithm. Additionally, a detailed analysis comparing the resulting maps with ground-truth Lidar-based maps is provided to highlight the algorithm’s accuracy and reliability.
@@ I Introduction
Simultaneous localization and mapping (SLAM) is a an essential process, involving self-localizing and construction of a map of the environment, especially important for self-navigating autonomous robots. Current state-of-the-art approaches rely heavily on sensors like cameras, IMUs (Inertial Measurement Units), laser range finders, and ultrasound. These sensors provide data that is used to generate accurate estimates of the environment map. However, cameras, laser range finders, and ultrasound sensors can be obstructed, corrupted by noise, or cause privacy concerns. Therefore, alternative approaches are essential when these exteroceptive sensors are unreliable or insufficient. Developing robust methods that can compensate for these limitations is critical to improving overall performance and ensuring accurate environment mapping in diverse and challenging conditions.
@@
As WiFi networks occupy many indoor and public spaces, the use-case of information extruding from WiFi systems is a promising alternative for achieving SLAM accuracy. Such information is known as received signal strength indicator (RSSI) data, protruding outward from WiFi emitters. This method is particularly advantageous in situations where traditional sensors such as cameras, lasers, and ultrasound may be inadequate, such as in environments with privacy concerns or poor lighting conditions [1]. The recent progress in this domain, such as WiFiSLAM[2], has predominantly concentrated on WiFi-based localization and the potential estimation of WiFi router positions, without generating a geometric map of the environment, needed for many robotic applications. There have been works relying on crowdsourcing [3], showing that it is possible to generate an environment map using WiFi data. However, these methods often face challenges such as limited accuracy due to signal interference, variability in signal strength, and the need for extensive data from multiple users for reliable results. Additionally, processing of the map with primarily RSSI information in the algorithms of many works, shows that it fluctuates unstably, being unreliable, particularly in environments that are dynamically changing [4]. Thus, it is clear that there exists a gap in the development of a way in which to leverage a WiFi-based mapping technique, from geometric constraints alone. The main problem arises in acquiring, from the WiFi data, an accurate geometric shape of the true mapped environment.
@@
This paper [^1] This paper is partly based on our earlier paper [5], extended with multiple routers, metric evaluations, and more experiments. presents a novel technique capable of generating geometric maps from WiFi signals. The method utilizes developments in computational geometry, namely the $k$-visibility algorithm [6]. We extend the $k$-visibility by presenting the concept of inverse $k$-visibility. This novel concept is then utilized to develop a geometric map of an unknown environment from WiFi signals, coined Structure from WiFi (SfW).
@@
Knowing accurately the mapped free space is imperative for the control of many autonomous systems, which need to know how to plan paths, making sure not to collide with occupied and obstacle spaces. The contributions of our work are: [^2] https://sites.google.com/view/structure-from-wifi/home 1) A novel algorithm that is capable of generating geometric maps using WiFi signals received from multiple routers, 2) benchmarking the WiFi-generated maps with Lidar-generated maps by comparing the area, number of data points, $k$-value prediction True/False setting, $k$-value accuracy percentage, IOU and MSE scores, and 3) evaluation on real-world collected from indoor spaces.
@@
Figure 1: The concept of $k$-visibility is demonstrated where different $k$-values are shown: $k=0$ (red), $k=1$ (green), $k=2$ (blue) and $k=3$ (yellow). Based on a straight-line measure from a reference point such as a router (shown in dark blue) to a desired location, $k$-visibility is a metric on how many times traverses through a wall/obstacle.
@@
The rest of this work is organized as follows: Sec. II presents an overview of recent approaches to WiFiSLAM and challenges faced in applications. Sec. III describes the background of $k$-visibility algorithm. Sec. IV proposes the inverse $k$-visibility algorithm. Sec. V extends the inverse $k$-visibility algorithm to real-world experiments utilizing sparse information. Sec. VI presents the experimental results. Finally, Sec. Future work and Conclusions discusses future works, limitations, and concludes the paper.
@@ II Literature Review
WiFi-based localization techniques have seen extensive adoption in recent years. Ferris et al. [2] succeeded in localization by converting high-dimensional signal strength data into a two-dimensional latent space and resolving SLAM using Gaussian Process Latent Variable Modeling (GP-LVM). This work was followed by other studies, including, [7, 8]. Graph-based algorithms for SLAM have also been explored, as in  [9, 10, 11, 4]. Enhancements in WiFi observation models have shown potential for high-accuracy mapping, as described in the research by [1, 12]. Both learning and non-learning methods have their advantages and disadvantages, depending on the robotic system. We will first address learning-based WiFi map construction approaches (See Sec. II-A), and go through non-learning, geometric WiFi-map methods (See Sec. II-B)
II-A Learning-Based WiFi Approaches
@@
The prominence of learning-based WiFi map construction has alleviated many issues with conventional geometric methods, but with the disadvantage of increased computational need in its pre- and post-processing steps  [13], [14]. Zou et al.  [15] introduced WiGAN, a generative adversarial network (GPR-GAN) for constructing detailed indoor radio maps using Gaussian process regression (GPR) with received signal strengths (RSS). Other notable works include  [16],  [17],  [18], and  [19]. Zhang et al.  [20] proposed a deep fuzzy forest to address the limitations of decision trees and deep neural networks in end-to-end training for WiFi-based indoor robot positioning. Additionally, Ayinla et al.  [21] developed SALLoc, a WiFi fingerprinting indoor localization scheme using Stacked Autoencoder (SAE) and Attention-based Long Short-Term Memory (ALSTM), which demonstrated high accuracy and robustness in large-scale indoor environments. Wang et al.  [22] introduced MapLoc, an LSTM-based indoor localization system that uses uncertainty maps created from magnetic field readings and WiFi RSS, showing superior performance in location prediction. Park et al.  [23] reviewed various machine learning techniques for WiFi-based indoor positioning, emphasizing their potential to enhance accuracy and scalability. Turgut and Gorgulu Kakisim  [24] proposed a hybrid deep learning architecture combining LSTM and CNN for WiFi-based indoor localization, achieving higher accuracy than baseline methods.
II-B Non-Learning-Based WiFi Approaches
@@
The alternative to Deep learning (DL) and Machine learning (ML) modelling deals with non-learning methods. Several studies have leveraged the sensors integrated into smartphones to infer indoor floor plans through a crowdsensing approach [3]. Research in this area often utilizes inertial sensors alongside WiFi signal strength data, both of which are common in commercial smartphones. Key papers include  [25, 26, 27, 28, 29, 30, 31, 32]. The main limitation of these studies is that they only produce “traversable maps”, identifying navigable areas while leaving occupied zones uncharted. Additionally, these works rely on crowdsourced or public data, which may not always be accessible, thereby restricting their applicability.
@@
In a different approach, Gonzalez-Ruiz and Mostofi [33] introduced a method to generate a non-invasive occupancy grid map using wireless measurements and directional antennas. Their framework employed a coordinated robot setup to create a 2D map of an environment by collecting data through walls and other obstaceles. Similar techniques have been applied in Ultra Wide-band (UWB) SLAM studies, utilizing UWB signal path propagation models with directional antennas to achieve similar goals [34, 35].
@@
Further, non-learning methods that do not rely on model predictions, such as Tong et al.  [36], achieve decimeter-level localization accuracy by segmenting Wi-Fi access points (APs) into groups. These methods generate local maps from the APs and then merge them to create a comprehensive global Wi-Fi map. Ninh et al.  [37] proposed a random statistical method for WiFi fingerprinting, achieving a maximum positioning error of less than 0.75 meters. Shi et al.  [38] introduced a training-free indoor localization approach using WiFi round-trip phase and factor graph optimization, achieving a mean absolute error (MAE) of 0.26 meters. Estrada et al.  [39] developed an indoor positioning system using OpenWRT, achieving an average margin error of 2.43 meters. Tao and Zhao  [40] presented algorithms leveraging extreme values for AP selection and positioning, demonstrating improved accuracy. Sulaiman et al.  [41] explored radio map generation methods to enhance RSSI-based indoor positioning, achieving approximately 0.45 meters accuracy. Azaddel et al.  [42] proposed SPOTTER, a WiFi and BLE fusion method based on particle filtering, showing a 35% improvement in accuracy. Tao et al.  [43] introduced CBWF, a lightweight WiFi fingerprinting system achieving an average localization accuracy of 2.95 meters. Tang et al.  [44] combined WiFi and vision for map construction and maintenance, enhancing accuracy and stability. Finally, Jurdi et al.  [45] presented WhereArtThou, a WiFi-RTT-based indoor positioning system, achieving 90th percentile distance errors of 1.65 meters, suitable for commercial deployment.
@@ III Background: $k$-visibility
The concept of $k$-visibility offers a unique perspective on WiFi systems and will be used through out our algorithms. This concept was initially introduced as the modem illumination problem [46, 47]. $k$-visibility investigates the extent of area that can be observable from the router points at the vertices of a polygon. This concept is an extension of the basic visibility problem, which focuses on determining the region that a single point within a polygon can see [6].
In a simple polygon P, two points p and q are considered to be mutually visible if the straight line connecting them does not intersect the exterior of P [6]. The visibility idea came from the Art Gallery Problem posed by Victor Klee in 1973 [48], which seeks to determine the minimum number of guards required to monitor an art gallery with n walls. The area that can be seen from a specific point forms what is known as a visibility polygon [6].
On the other hand, two points p and q are defined as $k$-visible if the line segment connecting them intersects the exterior of polygon P no more than k times [6]. Consequently, for a given k, it is possible to determine the $k$-visible region from a particular vantage point p.
Fig. 1 shows a map displaying various $k$-visibility values from the viewpoint of a point marked in dark blue within the red area. Here, red cells indicate 0-visibility, green cells represent 1-visibility, blue cells correspond to 2-visibility, and yellow cells denote 3-visibility.
The vantage point p is referred to as a k-transmitter, and $k$-visible polygons can be identified for each value in the set $\{0,1,..,k\}$ [6]. Several algorithms exist to generate $k$-visibility plots, each with varying levels of computational complexity [49, 50, 51].
@@ IV Dense Inverse $k$-visibility
It this section, we will go over the preliminaries and concept of the proposed dense inverse $k$-visibility, starting with an overview of the algorithm involved, with focus in the next section its application in robotics.
$k$-visibility deals with the numeric assignment of a region of space, based on the visibility of that space in reference to some known $k$-router or $k$-transmitter. It is based on ray casting in various directions, with many $k$-visibility works casting from $k$-router-point to current trajectory point, with $k$-value dependent on whether the casted ray intersects with an obstruction or obstacle. The rule is in $k$-visibility, the ray is made of subparts, with different $k$-values at these obstructions. These points along the ray where $k$ changes are namely where walls or occupied space is detected. A wall is located at the coincident point between two adjacent differing $k$-values.
The inverse $k$-visibility algorithm is based in the reverse of the $k$-visibility concept, where from the $k$-value regions, instead of using the the environment map as input, these $k$-value regions help to provide the map as output instead. This algorithm does so by doing ray-casting and deeming the changes between $k$-values as a wall location or occupied space. Fig. 2 demonstrates this ray drawing principle.
In this regard, if the entire set of points representing the change in $k$-values is found, then the full information of the environment map can be extracted by this inverse $k$-visibility algorithm. Color-coded according to differing $k$-values can be done and this can be used as input code map of all superimposed $k$-visibility plots for all values $k$, with respect to a known $k$-router or $k$-transmitter reference position, where in which the entire $k$-visibility is based on, as in Fig. 1. It is clear that wall and occupied positions lie on the intersection between consecutive $k_{i-1}k_{i}$ portions of the $k$-visible map. The details of the full inverse $k$-visibility algorithm are that it detects shared pixel spaces among successive $k$-value areas, superposes them on a separate image, using black-and-white outlining to generate the environment map.

### v2

@@ Abstract
In recent years, the increased availability of WiFi in indoor environments has gained interest in the robotics community to utilize WiFi signals for indoor simultaneous localization and mapping algorithms. This paper discusses the challenges of achieving high-accuracy geometric map building using WiFi signals. The paper introduces the concept of inverse k-visibility, developed from the k-visibility algorithm, to identify free space in an unknown environment, used for planning, navigation, and obstacle avoidance. Comprehensive experiments, including those utilizing single and multiple RSSI signals, were conducted in both simulated and real-world environments to demonstrate the robustness of the proposed algorithm. Additionally, a detailed analysis comparing the resulting maps with ground-truth LiDAR-based maps is provided to highlight the algorithm’s accuracy and reliability.
@@ I Introduction
Simultaneous localization and mapping (SLAM) is an essential process, involving self-localizing and the construction of a map of the environment, especially important for autonomous robots. Current state-of-the-art approaches rely heavily on sensors like cameras, inertial measurement units (IMUs), laser range finders, and ultrasound. These sensors provide data that is used to generate accurate estimates of the environment map. However, cameras, laser range finders, and ultrasound sensors can be obstructed, corrupted by noise, or cause privacy concerns. Therefore, alternative approaches are essential when these exteroceptive sensors are unreliable or insufficient. Developing robust methods that can compensate for these limitations is critical to improving overall performance and ensuring accurate environment mapping in diverse and challenging conditions.
@@
As WiFi networks occupy many indoor and public spaces, the use of information extracted from WiFi signals is a promising alternative for SLAM. One such information is known as the received signal strength indicator (RSSI), indicating the strength of the signal received by a WiFi receiver. RSSI is a function of the geometry of the environment and the location of the WiFi emitter and receiver. WiFi is particularly advantageous in situations where traditional sensors such as cameras, lasers, and ultrasound may be inadequate, such as in environments with privacy concerns or poor lighting conditions [kudo2017utilizing]. Methods such as WiFiSLAM [ferris2007wifi] have predominantly concentrated on WiFi-based localization and the potential estimation of WiFi router positions, without generating a geometric map of the environment, needed for many robotic applications. There have been works relying on crowdsourcing [zhou2021crowdsourcing], showing that it is possible to generate an environment map using WiFi data. However, these methods often face challenges such as limited accuracy due to signal interference, variability in signal strength, and the need for extensive data from multiple users for reliable results. Additionally, processing of the map with primarily RSSI information in the algorithms of many works shows that it fluctuates unstably, being unreliable, particularly in environments that are dynamically changing [arun2022p2slam]. Thus, a gap exists in developing a reliable WiFi-based mapping technique. The main problem arises in acquiring, from the WiFi data, an accurate geometric shape of the environment.
@@
This paper [^1] This paper is partly based on our earlier paper [kim2024structure], extended with multiple routers, metric evaluations, and more experiments. presents a novel technique capable of generating geometric maps from WiFi signals. The method utilizes developments in computational geometry, namely the $k$-visibility algorithm [o1987art], see Fig. 1. We extend the $k$-visibility by presenting the concept of inverse $k$-visibility. This novel concept is then utilized to develop a geometric map of an unknown environment from WiFi signals, coined Structure from WiFi (SfW).
@@
Knowing accurately the mapped free space is imperative for the control of many autonomous systems, which need to know how to plan paths, making sure not to collide with occupied and obstacle spaces or to minimize the number of collisions. The contributions of our work are: [^2] https://sites.google.com/view/structure-from-wifi/journal-extension 1) A novel algorithm that is capable of generating geometric maps using WiFi signals received from multiple routers, 2) benchmarking the WiFi-generated maps with LiDAR-generated maps by comparing the area, number of data points, $k$-value prediction True/False setting, $k$-value accuracy percentage, IOU and MSE scores, and 3) evaluation on real-world collected from indoor spaces.
@@
Figure 1: The concept of $k$-visibility is demonstrated where different $k$-values are shown: $k=0$ (red), $k=1$ (green), $k=2$ (black) and $k=3$ (yellow). Based on a straight-line measure from a reference point such as a router (shown in dark black) to a desired location, $k$-visibility is a metric on how many times the line traverses through a wall/obstacle.
@@
The rest of this work is organized as follows: Sec. II presents an overview of recent approaches to WiFi-based SLAM and challenges faced in applications. Sec. III describes the background of the $k$-visibility algorithm. Sec. IV proposes the inverse $k$-visibility algorithm. Sec. V extends the inverse $k$-visibility algorithm to real-world experiments utilizing sparse information. Sec. VI presents the experimental results. Sec. VII discusses current challenges and limitations of the work. Finally, Sec. VIII discusses future works and concludes the paper.
@@ II Literature Review
WiFi-based localization techniques have seen extensive adoption in recent years. Ferris et al. [ferris2007wifi] succeeded in localization by converting high-dimensional signal strength data into a two-dimensional latent space and resolving SLAM using Gaussian Process Latent Variable Modeling (GP-LVM). This work was followed by other studies, including, [xiong2017diversified, miyagusuku2016improving]. Graph-based algorithms for SLAM have also been explored, as in  [huang2011efficient, liu2019collaborative, herranz2016wifi, arun2022p2slam]. Enhancements in WiFi observation models have shown potential for high-accuracy mapping, as described in the research by [kudo2017utilizing, he2015wi]. Both learning and non-learning methods have their advantages and disadvantages, depending on the robotic system. We will first address learning-based WiFi map construction approaches (See Sec. II-A), and go through non-learning, geometric WiFi-map methods (See Sec. II-B)
II-A Learning-Based WiFi Methods
@@
The prominence of learning-based WiFi map construction has alleviated many issues with conventional geometric methods, but with the disadvantage of increased computational need in its pre- and post-processing steps  [Bellavista1], [LiuFen1]. Zou et al.  [zou1] introduced WiGAN, a generative adversarial network (GPR-GAN) for constructing detailed indoor radio maps using Gaussian process regression (GPR) with received signal strengths (RSS). Other notable works include  [Ayyalasomayajula1],  [Xuhen1],  [Karmanov1], and  [chenXi1]. Zhang et al.  [ZhangLe1] proposed a deep fuzzy forest to address the limitations of decision trees and deep neural networks in end-to-end training for WiFi-based indoor robot positioning. Additionally, Ayinla et al.  [Ayinla1] developed SALLoc, a WiFi fingerprinting indoor localization scheme using Stacked Autoencoder (SAE) and Attention-based Long Short-Term Memory (ALSTM), which demonstrated high accuracy and robustness in large-scale indoor environments. Wang et al.  [Wang1] introduced MapLoc, an LSTM-based indoor localization system that uses uncertainty maps created from magnetic field readings and WiFi RSS, showing superior performance in location prediction. Park et al.  [Park1] reviewed various machine learning techniques for WiFi-based indoor positioning, emphasizing their potential to enhance accuracy and scalability. Turgut and Gorgulu Kakisim  [Turgut1] proposed a hybrid deep learning architecture combining LSTM and CNN for WiFi-based indoor localization, achieving higher accuracy than baseline methods.
II-B Non-Learning WiFi Methods
@@
The alternative to deep learning (DL) and machine learning (ML) modelling deals with non-learning methods. Several studies have used the sensors integrated into smartphones to infer indoor floor plans through a crowdsensing approach [zhou2021crowdsourcing]. Research in this area often utilizes inertial sensors alongside WiFi signal strength data, both of which are common in commercial smartphones. Key papers include  [luo2014piloc, shin2011unsupervised, zhou2015alimc, zhou2018graph, shen2013walkie, jiang2013hallway, alzantot2012crowdinside, liang2016sensewit]. The main limitation of these studies is that they only produce “traversable maps”, identifying navigable areas while leaving occupied zones uncharted. Additionally, these works rely on crowdsourced or public data, which may not always be accessible, thereby restricting their applicability.
@@
In a different approach, Gonzalez-Ruiz and Mostofi [gonzalez2013cooperative] introduced a method to generate a non-invasive occupancy grid map using wireless measurements and directional antennas. Their framework employed a coordinated robot setup to create a 2D map of an environment by collecting data through walls and other obstacles. Similar techniques have been applied in Ultra Wideband (UWB) SLAM studies, utilizing UWB signal path propagation models with directional antennas to achieve similar goals [deissler2010uwb, deissler2012infrastructureless].
@@
Further, non-learning methods that do not rely on model predictions, such as Tong et al.  [tong1], achieve decimeter-level localization accuracy by segmenting Wi-Fi access points (APs) into groups. These methods generate local maps from the APs and then merge them to create a comprehensive global Wi-Fi map. Ninh et al.  [Ninh1] proposed a random statistical method for WiFi fingerprinting, achieving a maximum positioning error of less than 0.75 meters. Shi et al.  [Shi1] introduced a training-free indoor localization approach using WiFi round-trip phase and factor graph optimization, achieving a mean absolute error (MAE) of 0.26 meters. Estrada et al.  [Estrada1] developed an indoor positioning system using OpenWRT, achieving an average margin error of 2.43 meters. Tao and Zhao  [Tao1] presented algorithms utilizing extreme values for AP selection and positioning, demonstrating improved accuracy. Sulaiman et al.  [Sulaiman1] explored radio map generation methods to enhance RSSI-based indoor positioning, achieving approximately 0.45 meters accuracy. Azaddel et al.  [Azaddel1] proposed SPOTTER, a WiFi and BLE fusion method based on particle filtering, showing a 35% improvement in accuracy. Tao et al.  [TaoYe1] introduced CBWF, a lightweight WiFi fingerprinting system achieving an average localization accuracy of 2.95 meters. Tang et al.  [Tang1] combined WiFi and vision for map construction and maintenance, enhancing accuracy and stability. Finally, Jurdi et al.  [Jurdi1] presented WhereArtThou, a WiFi-RTT-based indoor positioning system, achieving 90th percentile distance errors of 1.65 meters, suitable for commercial deployment.
@@ III Background: $k$-visibility
The concept of $k$-visibility offers a unique perspective on WiFi systems and will be used throughout our algorithms. This concept was initially introduced as the modem illumination problem [fabila2009modem, aichholzermodem]. $k$-visibility investigates the extent of the area that can be observed from the router points at the vertices of a polygon. This concept is an extension of the basic visibility problem, which focuses on determining the region that a single point within a polygon can see [o1987art]. In a simple polygon P, two points p and q are considered to be mutually visible if the straight line connecting them does not intersect the exterior of P [o1987art]. The visibility idea came from the Art Gallery Problem posed by Victor Klee in 1973 [klee1969every], which seeks to determine the minimum number of guards required to monitor an art gallery with n walls. The area that can be seen from a specific point forms what is known as a visibility polygon [o1987art].
On the other hand, two points p and q are defined as $k$-visible if the line segment connecting them intersects the exterior of polygon P no more than k times [o1987art]. Consequently, for a given k, it is possible to determine the $k$-visible region from a particular vantage point p.
Fig. 1 shows a map displaying various $k$-visibility values from the viewpoint of a point marked in dark black within the red area. Here, red cells indicate 0-visibility, green cells represent 1-visibility, black cells correspond to 2-visibility, and yellow cells denote 3-visibility.
The vantage point p is referred to as a k-transmitter, and $k$-visible polygons can be identified for each value in the set $\{0,1,..,k\}$ [o1987art]. Several algorithms exist to generate $k$-visibility plots, each with varying levels of computational complexity [bajuelos2012hybrid, bahoo2020computing, bahoo2019time].
@@
Algorithm 1 Dense Inverse k-Visibility Algorithm
1:Inputs:
2:  Reference point $R$ (e.g., WiFi router position)
3:  $k$-value of each point along the rays
4:Outputs:
5:  A map $M$ that classifies regions as free space or walls
6:Initialize:
7: Set the map $M$ to unknown values ($M_{ij}\leftarrow 127$ $\forall$ $i,j$)
8:for all rays $r\in\mathcal{R}$ cast from the reference point $R$ do
9:  for all points $p$ along ray $r$ do
10:   get $k\text{-val}_{p}$
11:   if $k\text{-val}_{p}=0$ then
12:    Mark the region between $R$ and $p$ as free in $M$
13:   else if $k\text{-val}_{p}\geq 1$ then
14:    Mark the point where $k$ changes as a wall in $M$
15:   end if
16:  end for
17:end for
18: Return the final map $M$
@@ IV Dense Inverse $k$-visibility
In this section, we will go over the preliminaries and concept of the proposed dense inverse $k$-visibility, starting with an overview of the algorithm involved.
The $k$-visibility problem, explained in the previous sections, takes the location of a router and the geometry of the environment, e.g., the walls, as inputs. It then generates a value for each point in the space, which indicates how many walls exist between the point of the router. This value, known as the $k$-value, is calculated from RSSI in this paper. The inverse of this problem is of interest, where we assume that for each point in the space, we know its $k$-value as well as the location of the router. The objective is to determine the geometry of the environment, e.g., the walls. This is useful for making a geometric map of the environment for robotic applications. Since we assume the $k$-value is known for ‘all’ points, the inverse problem is coined the ‘dense’ inverse $k$-visibility. Mathematically, this can be written as:
  • $k$-visibility
    $\color[rgb]{0,0,0}\text{k\text{-}val}_{p}=\text{k\_vis}(R,p\mid map),\ \forall\,p\in S$}, (1)
  • dense inverse $k$-visibility
    $\color[rgb]{0,0,0}map(p)=\text{dense\_inv\_k\_vis}(R,\text{kval}_{p},p),\ \forall\,p\in S$}, (2)
where $S$ is the desired space (which is 2D in this paper), and $k\text{-}val_{p}$ is the $k$-value for point $p\in S$. $R$ is the 2D location of the router, and mapmap is the geometry (2D map) of the environment, represented as an occupancy grid map. We write $\text{k\_vis}(R,p\mid map)$ to indicate evaluation with respect to the geometry.
If the entire set of points representing the change in $k$-values is found, the geometry of the environment map can be extracted by the inverse $k$-visibility algorithm. As shown in Fig. 1, the wall and occupied cells lie on the intersection between consecutive $(k_{i-1},k_{i})$ portions of the $k$-visible map. The $k$-visibility maps must be aligned first such that the location of the router is the same in all of them.
Alg. 1 shows the details of the ‘dense inverse $k$-visibility’ algorithm. It works by reversing the traditional $k$-visibility approach, where $k$-value regions are used to generate a map of the environment. The inputs are the 2D position of the router and $k$-values of each point along the ray coming out of the router (lines 2, 3). The output is a 2D occupancy map showing the free and occupied cells (line 5). Every cell of the map is initialized as unknown (line 7). The algorithm begins by casting rays $\mathcal{R}$ from a reference point $R$ (line 8), and for each point $p$ along the ray (line 9), the $k$-value is retrieved. The values are calculated from RSSI values, which will be explained in the next section. If the $k$-value at a point $p$ is zero (line 11), meaning no obstacles are encountered, the algorithm marks the entire region between $R$ and $p$ as free space or 255 (line 12). On the other hand, if $k(p)\geq 1$ (line 13), the algorithm identifies the point where the $k$-value changes and marks this as the location of an obstacle or 0 (line 14). The algorithm iterates through all rays $r\in\mathcal{R}$ (line 8), progressively updating the map $M$. The final map $M$ is returned once all rays have been processed, providing a detailed map of the environment. The function that maps signal strength to $k$-values will be discussed in detail in Sec. V.